Border Steel invested $500,000 in a new shearing unit. At an interest rate of 12% per year compounded quarterly, the annual income required to recover the investment in 3 years is closest to:

(a) $283,552

(b)$226,235

(c)$236,237

(d)$210,148
Answer #1

Ans.-(D)

Future value of 500000 after 3 years at interest rate of 12% per year compounded quarterly = 500000(1+0.03)^12 = 712880.44. ( Quarterly interest = 12%/4 = 3% and there are 12 quarters in 3 years)

If 210148 is received at the end of each year it's future value is

210148(1+0.03)^8+210148(1.03)^4+210148 = 712880.62